OpenStudy (anonymous):

Determine the infinite limit? (I have to present it in class. can I have an explanation please). lim + x + 2 x->-3 ------- = x + 3

OpenStudy (anonymous):

+ x approach -3

OpenStudy (anonymous):

\[ \lim_{x\to -3^+}\frac{x+2}{x+3}= \]

OpenStudy (zzr0ck3r):

@micalg do you understant?

OpenStudy (zzr0ck3r):

well as we aprroac-3 from the right we will have negative on top, and positive on bottom, but the bottom will tend to 0 as we get close so we have -/+ but the top is a constant, so limit = -infinity

OpenStudy (anonymous):

Fill out a chart: 

OpenStudy (anonymous):

yah I think I started to

OpenStudy (anonymous):

Also you can use epsilon delta to prove it: \[ \forall N, \exists \delta: \\ \forall x \quad 0<x-(-3)<\delta \implies \frac{x+2}{x+3}<N \]

OpenStudy (zzr0ck3r):

it does not converge @wio so you would need\[\forall \space M,\exists\epsilon>0\space st\space |x-(-3)|<\delta \implies f(x)>M\]

OpenStudy (zzr0ck3r):

and its not N, unless you are saying that N is in R, but we never say that...

OpenStudy (anonymous):

This is a one sided infinite limit. It has a particular definition. I might have made a minor error.

OpenStudy (zzr0ck3r):

if it has a limit, this does not.

OpenStudy (anonymous):

But there should not be any absolute value signs for this type of limit.

OpenStudy (zzr0ck3r):

evrey definition I know of limits has abs value...

OpenStudy (anonymous):

The one-sided limit does exist in this case.

OpenStudy (anonymous):

There are about 12 variations on the definition of a limit to account for things like one-sided limits, infinite limits, limits of sequences, etc.

OpenStudy (anonymous):

Actually much more than 12 variations. Anyway the general definition takes a lot of math that is not introduced in calculus to write out well.

OpenStudy (zzr0ck3r):

its not that bad, I was simply saying that you are forcings things to be smaller than a natural number, it should be all real numbers

OpenStudy (anonymous):

I didn't mean to say anything about natural or real numbers.

OpenStudy (zzr0ck3r):

so it would be \[\forall M , \exists\delta>0, st\space -\delta<x<a\implies f(x)>m\]

OpenStudy (zzr0ck3r):

we are showing divergence so f(x)<N makes no sense...

OpenStudy (anonymous):

I intially said \(f(x) < N \) is because before you said \(-\infty\).

OpenStudy (anonymous):

But since it is going towards \(\infty\) then yes, in this case you'd do \(f(x)>N\)